# Predicting post-release defects with knowledge units (KUs) of programming languages: an empirical study

### v1

@@ 5.1.2 Approach
    Here, 1 represents the highest possible AUC score. For example, to calculate the normalized AUC improvement of KUCLS over CC_PROD, where the AUC of KUCLS is 0.81 and the AUC of CC_PROD is 0.75, the calculation is as follows: the normalized AUC improvement is (0.81-0.75) / (1 - 0.75) $\times$ 100 % = 24%. This result indicates that KUCLS’s performance is 24% closer to the maximum possible improvement compared to the baseline. Since we construct 100 bootstrap models, we compute the average normalized AUC improvement across these models to provide a comprehensive measure of overall performance improvement.
Table 4: P-value, effect size, and median AUC value of KUCLS, CC_PROD, and CC defect prediction models.
Statistical Test
Median AUC KUCLS vs CC_PROD KUCLS vs CC
Release KUCLS CC_PROD CC P-value Effect Size P-value Effect Size
activemq-5.0.0 0.82 0.81 0.89 <0.05 M(0.44) <0.05 L(-1.00)
activemq-5.1.0 0.83 0.80 0.82 <0.05 L(0.59) <0.05 N(0.12)
activemq-5.2.0 0.83 0.80 0.85 <0.05 M(0.46) <0.05 L(-0.59)
activemq-5.3.0 0.82 0.79 0.83 <0.05 M(0.40) <0.05 M(-0.40)
activemq-5.8.0 0.85 0.84 0.86 <0.05 S(0.26) <0.05 M(-0.44)
derby-10.2.1.6 0.74 0.77 0.85 <0.05 L(-0.87) <0.05 L(-1.00)
derby-10.3.1.4 0.84 0.85 0.85 <0.05 S(-0.32) <0.05 L(-0.49)
derby-10.5.1.1 0.82 0.82 0.84 >0.05 N(0.01) <0.05 L(-0.62)
groovy-1.5.7 0.82 0.78 0.79 <0.05 S(0.29) <0.05 S(0.20)
groovy-1.6.BETA_1 0.82 0.81 0.87 <0.05 S(0.18) <0.05 L(-0.60)
groovy-1.6.BETA_2 0.85 0.83 0.91 <0.05 L(0.61) <0.05 L(-0.77)
hbase-0.94.0 0.72 0.78 0.84 <0.05 L(-0.95) <0.05 L(-1.00)
hbase-0.95.0 0.72 0.76 0.81 <0.05 L(-0.93) <0.05 L(-1.00)
hbase-0.95.2 0.71 0.69 0.72 <0.05 S(0.19) <0.05 L(-0.87)
hive-0.10.0 0.72 0.78 0.83 <0.05 L(-0.96) <0.05 L(-1.00)
hive-0.12.0 0.80 0.72 0.84 <0.05 L(0.99) <0.05 L(-1.00)
hive-0.9.0 0.79 0.82 0.90 <0.05 L(-0.96) <0.05 L(-1.00)
jruby-1.1 0.86 0.82 0.85 <0.05 L(0.54) <0.05 S(0.15)
jruby-1.4.0 0.83 0.81 0.84 <0.05 S(0.17) <0.05 N(-0.13)
jruby-1.5.0 0.84 0.83 0.86 <0.05 S(0.27) <0.05 S(-0.29)
jruby-1.7.0.preview1 0.84 0.81 0.83 <0.05 L(0.56) <0.05 S(0.17)
lucene-2.3.0 0.85 0.82 0.91 <0.05 L(0.65) <0.05 L(-1.00)
lucene-2.9.0 0.79 0.78 0.82 <0.05 S(0.19) <0.05 L(-0.8)
lucene-3.0.0 0.91 0.89 0.9 <0.05 L(0.60) <0.05 S(0.33)
lucene-3.1.0 0.74 0.72 0.76 <0.05 M(0.39) <0.05 L(-0.71)
wicket-1.3.0-beta2 0.81 0.80 0.82 <0.05 S(0.21) <0.05 S(-0.26)
wicket-1.3.0-incubating-beta-1 0.83 0.84 0.87 >0.05 N(-0.07) <0.05 L(-0.61)
wicket-1.5.3 0.81 0.77 0.79 <0.05 L(0.61) <0.05 M(0.43)
Median 0.82 0.80 0.84
@@ 5.1.3 Findings
Observation 4) The minimum AUC of the KUCLS is above 0.70 (i.e., AUC > 0.70), indicating that KUs are a good candidate to preeict post-release defects. Table 4 shows the P-value, effect size, and median AUC value of the prediction models for each studied release. The KUCLS column refers to the AUC of the models built with KU-based features. We observe that the AUCs of the KUCLS are always above 0.70 (minimum is 0.72). In particular, the median AUC of the KUCLS across all the studied releases is 0.82. A defect prediction model is typically considered for use in practice when its AUC is higher than or equal to 0.70 (Jiarpakdee et al., 2020). This result highlights that KUs are a good candidate to classify post-release defects.
Observation 5) The KUCLS outperforms the CC_PROD in classifying post-release defects. As we observe in Table 4, the median of the CC_PROD column is 0.80, whereas the median of the KUCLS column is 0.82. We also compare the individual releases, and we observe that the KUCLS outperforms the CC_PROD in the majority of the studied releases. We observe that the KUCLS outperforms the CC_PROD in 20 of the 28 releases (see green rows in Table 4 of the column named “KUCLS vs CC_PROD”), the majority of them with a large and medium effect size. The effect size is negligible in 2 out of 28 releases, indicating that both models perform the same in those cases. However, for the remaining 6 releases, the KUCLS does not outperform the CC_PROD, but maintain a good AUC value (the median AUC of KU models for these 6 releases is 0.73). In summary, the KUCLS wins in 20 releases, the CC_PROD wins in 6 releases, and a draw occurs in the remaining cases. We also calculate the normalized AUC improvement between the KUCLS and the CC_PROD. The normalized AUC improvement of the KUCLS over the CC_PROD is between 5.1% and 28.9% across the studied releases. These results show that the KUCLS outperform the CC_PROD.
Observation 6) The KUCLS generally does not outperform the CC in most of the studied releases. However, the KUCLS still achieves strong AUC values. As we observe in Table 4 the median of the CC column is 0.84, whereas the median of the KUCLS column is 0.82. We observe that the KUCLS outperforms the CC in 5 of the 28 studied releases. For the remaining releases (23 of the 28 studied releases), CC models outperform KUCLS models. The median AUC values of KUCLS for those 23 releases is 0.82. Therefore, while the KUCLS does not outperform the CC, the KUCLS still achieves strong AUC values.
Summary RQ1: How effective are KU-based features for identifying defect prone code? KUs are good candidates for predicting bug-prone code. In particular: The KUCLS can classify post-release defects with a minimum AUC of 0.73 and a median AUC of 0.82. The KUCLS outperforms the CC_PROD in the defect prediction task. The normalized AUC improvement of the KUCLS over the CC_PROD ranges from 5.1% to 28.9% across the studied releases. While the KUCLS does not outperform the CC, it maintains strong AUC values.
5.2 RQ2: What are the important KU-based features of the KUCLS for predicting post-release defects?
5.2.1 Motivation
Studying the feature importance of the KUCLS is essential to identify which KU-based feature contribute most significantly to predicting post-release defects. This analysis not only enhances our understanding of the model’s decision-making process but also highlights the practical relevance of specific KUs. From an operational perspective, identifying key features streamline the feature engineering process and improves the model’s practicality for deployment in real-world scenarios.
5.2.2 Approach
To determine the KU-based features that have the most significant impact on predictions, we conduct a model interpretation analysis utilizing SHAP (SHapley Additive exPlanation). SHAP, a widely adopted method for model interpretation, leverages game theory to provide an optimal framework for quantifying each feature’s contribution to the model’s predictions (Molnar, 2020). SHAP values can be positive or negative, signifying the direction of the feature’s influence on the prediction. A positive SHAP value indicates that the feature increases the likelihood of the predicted outcome (e.g., classifying a file as buggy in a defect prediction model). In contrast, a negative SHAP value means that the feature reduces the likelihood of the predicted outcome or increases the likelihood of the opposite class (i.e., not buggy class). The strength of a feature is determined by the magnitude of SHAP values. Specifically, the absolute SHAP value of a feature indicates its influence on the prediction, with higher values signifying greater importance. We use the SHAP package of python to perform the SHAP analysis (CRAN, 2020).
We compute KU-based feature importance for the KUCLS as follows. First, we apply SHAP to the KUCLS and calculate the absolute SHAP value for each feature across every file (i.e., record) in the dataset of a release. Next, for each feature, we sum the SHAP values of all files (i.e., records) in the dataset. We follow this approach for all the 100 bootstraps of every release. As a result we obtain a distribution of the sum of SHAP values for every feature across all releases. Then, we apply the Scott-Knott ESD technique on the distribution of the sum of SHAP values generated from the 100 bootstraps of every release. As a result, we obtain a feature ranking for each release. For example, in release jruby-1.1, the most important feature of the KUCLS is Inheritance (rank 1), the second most important feature is Exception (rank 2) and so on. There can be ties, i.e., two features with same importance are given the same rank. By taking all releases, we can produce a feature rank distribution for each feature. To summarize the results across all releases, we apply the Scott-Knott ESD technique again on the feature rank distributions, similarly to Ghotra et al. (2015). The SK-ESD is a method of multiple comparison that leverages a hierarchical clustering to partition the set of treatment values (e.g., medians or means) into statistically distinct groups with non-negligible difference (Mittas and Angelis, 2013). As a result, we obtain a final rank for each feature. We use Tim Menzies’s implementation of the Scott-Knott ESD technique, which employs non-parametric statistical tests and the Cliff’s Delta measure of effect size (Menzies, 2020).
Figure 7: The rank distribution of features of the KUCLS for defect classification. The number inside the square brackets indicates the final rank of the feature after applying Scott-Knott ESD for the second time.
5.2.3 Findings
Observation 7) Method Encapsulation is the most important KU-based feature across all studied releases of the KUCLS for predicting post-release defects. Figure 7 shows the rank distribution of each feature (final rank is shown in brackets). We observe that Method Encapsulation is the top-ranked KU in the KUCLS. Indeed, the boxplot reveals that Method Encapsulation KU is the most important feature in most of our studied releases as the median value is one and the interquartile range is narrow (most ranks are one). Inheritance KU is ranked second and Exception KU is the third-ranked feature. We then have two KU-based features ranked forth, namely: Advanced Class Design and LOOP. Of all these features, we note that Advanced Class Design KU is the one with the most stable rank (smallest interquartile range). The other KU is a little less stable, meaning that its rank varies more across the studied releases. Some advanced KUs derived from Java SE II certification exams also rank among the top ten most important features. For instance, Generics and Collections and Java IO are both ranked eighth and exhibit stable rankings (small interquartile range). While Concurrency KU ranks eleventh, its ranking shows greater variability with a large interquartile range. KUs related to building enterprise-level applications in Java, such as Enterprise JavaBeans and Java Message Service (JMS), which are derived from Java EE certification exams, are generally ranked lower in importance, with most being ranked 14th. However, they also have significant contributions in predicting post-release defects. These advanced Java EE features are less frequently used in the studied systems compared to more common KU-based features like Data Type and Loop. As a result, they may have minimal influence on the global defect prediction model.
Summary RQ2: What are the important KU-based features of the KUCLS for predicting post-release defects? Method Encapsulation, Inheritance, and Exception are identified as the top three most influential KU-based features of the KUCLS for predicting post-release defects consistently across all studied releases.
5.3 RQ3: Can KU-based features of the KUCLS provide insights into the incidence of post-release defects?
5.3.1 Motivation
While global interpretations of defect prediction models provide a broad overview of the relationships between features and defects, they often fail to capture the nuances of specific instances. In contrast, instance-specific interpretations (a.k.a., local interpretations) offer detailed insights into the factors influencing predictions for individual source code (i.e., file). Such insights can reveal unique combinations of traits and contexts that contribute to defects, enabling a more targeted understanding of their occurrence. Hence, we investigate whether an instance-specific interpretation from the KUCLS can provide insights into the occurrence of defects or the context in which defects happen. We also compare this instance-specific interpretation with that given by the CC model. Our conjecture is that the two explanations can complement each other.
5.3.2 Approach
We select the activemq-5.8.0 release to study this research question. Our justification is that the median performance of the KUCLS and the CC are both above 0.85 for this release. To pinpoint and contrast the explanations given by each model, we perform a local interpretation analysis using SHAP. With SHAP, one can perform instance-specific (i.e., file level) interpretations. To select the instance, we first investigate the issues that are linked to the files. We parsed the commit messages of all commits that were submitted during the development of the activemq-5.8.0 release and extracted issues IDs from these messages. As a result, we established a mapping between defect-prone files in activemq-5.8.0 (as per our defect dataset) and issue IDs. We manually investigate a few defect-prone files and their corresponding issue tracking IDs in an ad-hoc manner and chose one case to showcase its results in this RQ. We choose the case where the priority of the issue was Major and the description of both the issue and the fix were easily understandable and clearly identifiable. Our chosen file was ActiveMQSession.java and the associated issue ID was AMQ-4634101010https://issues.apache.org/jira/browse/AMQ-4634. According to the issue reporter, the issue seems to be related to how ActiveMQ handles transactions in a Java Message Service (JMS) environment when working with distributed transactions.
Figure 8: The SHAP values of the top ten features of the CC and KUCLS models for their contribution to classifying the ActiveMQSession.java file as defect-prone.
5.3.3 Findings
Observation 8) KU-based features of the KUCLS offer insights into the characteristics of defect-prone files, which can complement the insights given by the CC model. Figure 8 presents the result of the SHAP analysis of the defective file for the CC and KUCLS models. As shown in Figure 8(a), the top-3 features that more strongly drive the CC model’s classification decision are: ADEV (i.e., the number of active developers), CountDeclMethodPublic (i.e., the number of local public methods) and CountClassCoupled (i.e., the number of other classes coupled to). The positive SHAP values of these code metrics highlight the positive contributions of the file to be classified as defect. Notably, the first two metrics, ADEV and CountDeclMethodPublic, have significantly higher positive SHAP values (both exceeding 0.10), compared to CountClassCoupled, which has a SHAP value of only 0.03. We do observe a high number of active developers (30), public methods (28) and coupled classes (32) in this class.
Figure 8(b) shows the KU-based features that drive the KUCLS model’s decision. The top three KU-based features that positively contribute to classifying the file as defect-prone are: Message Service, String Processing, and Java IO. Among these, Message Service has the highest positive SHAP value (0.14), significantly higher than the SHAP values of the other KU-based features. This result indicates that the Message Service KU is strongly associated with classifying this file as defect-prone. By manually investigating the patch [^11] https://shorturl.at/Kkc3L, we note that one of the changed methods (run(), line 860 does use Message Service APIs (an instantiation of an object of MessageAck class). In addition, the defect-prone class file instantiates and uses a set of methods and subclasses of several Message Service APIs (e.g., Session, Message, QueueSender and QueueReceiver). Therefore, it becomes clear that the defect-prone class (and the defect-prone method) do operate in a context where the concept of message service matters. Hence, we conclude that, in this case, KU-based features could provide further insights into the context in which the defect happened.
By combining the interpretations from code metrics and KU-based features, we now know two things: (i) the defect-prone class consists of too many methods and dependencies (ii) the defect-prone class and method operate in a context where message service matters. While these explanations might not necessarily lead to the discovery of causal relationships (which is a limitation that is actually inherent to the interpretation of machine learning models), we believe that they tell more about the context in which the defect happened and can thus inform future maintenance and refactoring decisions surrounding a given defect-prone class. We highlight that the instance discussed in this section is just one illustrative example of a concrete scenario. Nevertheless, the same approach (SHAP) could be used to investigate any defect-prone file from our dataset. Having said that, we encourage future work to investigate more defects and ask for feedback from the developers who worked on resolving these defects.
Summary RQ3: Can KU-based features of the KUCLS provide insights into the incidence of post-release defects? Yes. KU-based features offer insights into the characteristics of defect-prone files. Such insights are complementary to those offered by traditional code metrics.
5.4 RQ4: Can KUCLS be made more accurate by combining KU-based features with traditional code metrics?
5.4.1 Motivation
This research question investigates whether the predictive performance of the KUCLS for post-release bugs can be improved by integrating them with traditional code metrics. The underlying motivation is based on the hypothesis that combining diverse sets of features – KUs and traditional code metrics – can provide a more comprehensive understanding of the factors influencing post-release bugs. Such integration may enable the combined model to capture a wider range of predictive indicators, thereby enhancing its accuracy in identifying potential bugs.
5.4.2 Approach
We build a combined model named KUCLS+CC for predicting post-release bugs using all studied features of the KUCLS and all studied traditional code metrics of the CC (process and product metrics). Following the outlined steps in Step 1 of RQ1, we build the KUCLS+CC models for all studied releases. We compare the performance of the KUCLS+CC model with the performance of the KUCLS and CC models. We rank the studied prediction models (KUCLS, CC_PROD, CC and KUCLS+CC) using SK-ESD (Ghotra et al., 2015; Mittas and Angelis, 2013; Tantithamthavorn et al., 2017). We also analyze the feature importance of the KUCLS+CC using the same SHAP analysis method explained in Section 5.2.
5.4.3 Findings
@@
Figure 9: The distribution of AUC of KUCLS, CC, CC_PROD and the combined KUCLS+CC models. The models are grouped based on their performance rankings determined by the Scott-Knott ESD (SK-ESD) method, where a lower SK-ESD rank indicates a better-performing model.

### v2

@@ 4.1 RQ1: How effective are KU features for predicting post-release defects?
    Here, 1 represents the highest possible AUC score. For example, to calculate the normalized AUC improvement of a model M over baseline B, where the AUC of M is 0.81 and the AUC of B is 0.75, we would compute (0.81-0.75) / (1 - 0.75) $\times$ 100 % = 24%. This result indicates that M’s improvement is 24% of the maximum possible AUC improvement. Since we build 100 models for each studied system (i.e., one for each bootstrap sample), we report the average normalized AUC improvement when comparing two models.
    We also rank the studied prediction models (KUM , TM , PROD, PROC, OWN) using Scott-Knott ESD (SK-ESD) (Ghotra et al., 2015; Mittas and Angelis, 2013; Tantithamthavorn et al., 2017). The SK-ESD is a method of multiple comparison that leverages a hierarchical clustering to partition the set of treatment values (e.g., medians or means) into statistically distinct groups with non-negligible difference (Mittas and Angelis, 2013). We use Tim Menzies’ implementation of the SK-ESD technique, which employs non-parametric statistical tests and the Cliff’s Delta measure of effect size (Menzies, 2020).
Table 3: The effect size and median AUC value of KUM , PROD, PROC, OWN and TM defect prediction models.
Median AUC Statistical Test (Effect Size)
Release KUM PROD PROC OWN TM
KUM
vs PROD
KUM
vs PROC
KUM
vs OWN
KUM
vs TM
activemq-5.0.0 0.83 0.80 0.81 0.84 0.89 L(0.77)* L(0.63)* S(-0.28)* L(-1.00)*
activemq-5.1.0 0.83 0.80 0.62 0.73 0.82 L(0.76)* L(1.00)* L(0.99)* S(0.21)*
activemq-5.2.0 0.83 0.80 0.69 0.82 0.85 L(0.75)* L(1.00)* N(0.07) L(-0.70)*
activemq-5.3.0 0.82 0.80 0.66 0.81 0.84 L(0.64)* L(1.00)* S(0.29)* L(-0.58)*
activemq-5.8.0 0.85 0.83 0.68 0.74 0.86 L(0.61)* L(1.00)* L(1.00)* M(-0.41)*
derby-10.2.1.6 0.74 0.77 0.78 0.79 0.86 L(-0.73)* L(-0.95)* L(-0.99)* L(-1.00)*
derby-10.3.1.4 0.84 0.82 0.71 0.75 0.86 S(0.18)* L(1.00)* L(1.00)* L(-0.72)*
derby-10.5.1.1 0.83 0.82 0.60 0.75 0.85 M(0.44)* L(1.00)* L(1.00)* L(-0.87)*
groovy-1.5.7 0.81 0.77 0.53 0.71 0.78 M(0.38)* L(0.93)* L(0.68)* S(0.24)*
groovy-1.6.BETA_1 0.82 0.80 0.72 0.75 0.87 S(0.23)* L(0.79)* L(0.76)* L(-0.64)*
groovy-1.6.BETA_2 0.85 0.89 0.75 0.77 0.91 L(-0.5‘)* L(0.90)* L(0.83)* L(-0.64)*
hbase-0.94.0 0.71 0.78 0.64 0.76 0.83 L(-0.9‘)* L(0.97)* L(-0.78)* L(-1.00)*
hbase-0.95.0 0.71 0.69 0.67 0.70 0.82 L(0.90)* L(0.82)* S(0.22)* L(-1.00)*
hbase-0.95.2 0.71 0.65 0.63 0.69 0.73 L(0.69)* L(0.83)* M(-0.39)* L(-0.95)*
hive-0.10.0 0.72 0.78 0.67 0.79 0.85 L(-0.90)* M(0.43)* L(-0.93)* L(-1.00)*
hive-0.12.0 0.72 0.79 0.62 0.80 0.85 L(-0.97)* L(0.94)* L(-0.98)* L(-1.00)*
hive-0.9.0 0.79 0.86 0.57 0.70 0.91 L(-0.92)* L(1.00)* L(0.98)* L(-1.00)*
jruby-1.1 0.86 0.82 0.66 0.83 0.84 L(0.66)* L(1.00)* M(0.42)* M(0.49)*
jruby-1.4.0 0.83 0.82 0.67 0.73 0.85 M(0.45)* L(1.00)* L(0.98)* M(-0.44)*
jruby-1.5.0 0.84 0.82 0.75 0.82 0.86 M(0.45)* L(0.95)* M(0.39)* S(-0.28)*
jruby-1.7.0.preview1 0.84 0.81 0.51 0.76 0.85 L(0.74)* L(1.00)* L(0.90)* M(-0.39)*
lucene-2.3.0 0.82 0.84 0.88 0.85 0.92 M(-0.42)* L(-0.81)* L(-0.74)* L(-1.00)*
lucene-2.9.0 0.78 0.78 0.68 0.69 0.82 M(0.41)* L(1.00)* L(1.00)* L(-0.77)*
lucene-3.0.0 0.91 0.89 0.71 0.68 0.90 L(0.81)* L(1.00)* L(1.00)* M(0.46)*
lucene-3.1.0 0.73 0.72 0.64 0.61 0.75 S(0.27)* L(0.87)* L(0.96)* L(-0.56)*
wicket-1.3.0-beta2 0.82 0.80 0.74 0.70 0.83 M(0.44)* L(0.94)* L(1.00)* M(-0.43)*
wicket-1.3.0-incubating-beta-1 0.86 0.84 0.85 0.79 0.87 M(0.13)* M(0.38)* L(0.62)* L(-0.66)*
wicket-1.5.3 0.81 0.77 0.53 0.70 0.79 L(0.73)* L(1.00)* L(0.97)* M(0.39)*
Median 0.82 0.80 0.67 0.75 0.85
@@
Findings. Observation 1) The minimum AUC of KUM is above 0.70 (i.e., AUC > 0.70), indicating that KUs are a good candidate to predict post-release defects. Table 3 shows the effect size and median AUC value of the prediction models for each studied release. The KUM column refers to the AUC of the models built with the KU feature set. We observe that the AUCs of KUM are always above 0.70 (minimum is 0.71). In particular, the median AUC of KUM across all the studied releases is 0.82. A defect prediction model is typically considered suitable to be used in practice when its AUC is higher than or equal to 0.70 (Jiarpakdee et al., 2020). This result highlights that our KU features are good candidates to predict post-release defects.
@@
Figure 4: The distribution of AUC of KUM and other studied models. The models are grouped based on their performance rankings determined by the Scott-Knott ESD (SK-ESD) method, where a lower SK-ESD rank indicates a better-performing model.
Observation 2) KUM outperforms all baseline models except for TM. Figure 4 presents the SK-ESD ranks of the studied models across the studied releases. KUM ranks second, outperforming the baseline models that are built with one group of the traditional metrics (PROD, OWN and PROC). The median AUC values for these baseline models are 0.80 (PROD), 0.75 (OWN), and 0.67 (PROC), whereas, the median AUC for KUM is 0.82. Indeed, analysis of Table 3 indicates that KUM consistently outperforms these three baseline models in the majority of the studied releases. For example, KUM outperforms PROD in 21 out of the 28 studied releases, typically with either a medium or large effect size (note green rows in Table 3 of the column named “KUM vs PROD”). KU metrics can thus be interpreted as a set of product metrics with higher predictive power than traditional product metrics. Additionally, we calculate the normalized AUC improvement of KUM over PROD, PROC, and OWN for each release. The normalized AUC improvement of KUM over PROD ranges from 3.6% to 22.2% (median 11.4%), improvement over PROC ranges from 6.7% to 68.9% (median 42.4%), and improvement over OWN ranges from 3.3% to 71.9% (median 33.3%). In summary, KUM not only outperforms a model built on the same raw data (PROD), but also those built on process (PROC) and ownership (OWN) data.
Finally, while TM does outperform KUM in 23 out of the 28 releases (Table 3), KUM achieves a median AUC of 0.82 in those 23 releases. This result indicates that KUM would still be a robust classifier for those releases in practice.
Summary RQ1: How effective are KU features for predicting post-release defects? KUs are good candidates for predicting defect-prone code. In particular: KUM can classify post-release defects with a minimum AUC of 0.71 and a median AUC of 0.82. KUM outperforms the baseline models that are built with each individual group of traditional metrics (PROC, PROD and OWN). The normalized AUC improvement of KUM over PROD ranges from 3.6% to 22.2% (median 11.4%), improvement over PROC ranges from 6.7% to 68.9% (median 42.4%), and improvement over OWN ranges from 3.3% to 71.9% (median 33.3%). While KUM does not outperform TM (model built with all traditional metrics), KUM maintains strong AUC values.
4.2 RQ2: What are the most important KU features for predicting post-release defects?
Motivation. Analyzing feature importances helps pinpoint the features that contribute most significantly to predicting post-release defects. This analysis not only enhances our understanding of the model’s decision-making process but also highlights the practical relevance of specific KUs.
Approach. To determine the KU features that have the most significant impact on predictions, we conduct a model interpretation analysis utilizing SHAP (SHapley Additive exPlanation). SHAP is a widely adopted method for model interpretation that leverages game theory to provide an optimal framework for quantifying each feature’s contribution to the model’s predictions (Molnar, 2020). SHAP values can be positive or negative, signifying the direction of the feature’s influence on the prediction. A positive SHAP value indicates that the feature increases the likelihood of the predicted outcome (e.g., classifying a file as defective in a defect prediction model). In contrast, a negative SHAP value means that the feature reduces the likelihood of the predicted outcome or increases the likelihood of the opposite class (i.e., clean class). The strength of a feature is determined by the magnitude of SHAP values. Specifically, the absolute SHAP value of a feature indicates its influence on the prediction, with higher values signifying greater importance. We use the shapper python package to perform the SHAP analysis (CRAN, 2020).
We compute feature importances for KUM as follows. First, we apply SHAP to KUM and calculate the absolute SHAP value for each feature across every file (i.e., record) in the dataset of a release. Next, for each feature, we sum the SHAP values of all files (i.e., all records) in the dataset. We follow this approach for all the 100 bootstraps of every release. As a result we obtain a distribution of the sum of SHAP values for every feature across all releases. Then, we apply the SK-ESD technique (Mittas and Angelis, 2013) on the distribution of the sum of SHAP values generated from the 100 bootstraps of every release. As a result, we obtain a feature ranking for each release. Since there can be ties, two features with same importance are given the same rank. By taking all releases, we can produce a feature rank distribution for each feature. To summarize the results across all releases, we apply the SK-ESD technique again on the feature rank distributions, similarly to Ghotra et al. (2015). As a result, we obtain a final rank for each feature.
Figure 5: The rank distribution of features of KUM for defect classification. The number inside the square brackets indicates the final rank of the feature after applying Scott-Knott ESD for the second time.
Findings. Observation 3) Method & Encapsulation is the most important KU feature across all studied releases of KUM for predicting post-release defects. Figure 5 shows the rank distribution of each feature (final rank is shown in brackets). We observe that Method & Encapsulation is the top-ranked feature in KUM. Indeed, the boxplot reveals that Method & Encapsulation is the most important feature in most of our studied releases as the median value is one and the interquartile range is narrow (most ranks are one). Inheritance is ranked second and Exception Handling is the third-ranked feature. We then have two features ranked forth, namely: Advanced Class Design and Loop. Some advanced KUs derived from Java SE II certification exams also rank among the top ten most important features. For instance, Generics and Collections and Java IO are both ranked eighth. KUs related to building enterprise-level applications in Java, such as Enterprise JavaBeans and Java Message Service (JMS), which are derived from the Java EE certification exam, are ranked lower in importance (typically 14th). Upon closer inspection, we note that these advanced Java EE features are less frequently used (or even not used at all) in the studied releases. In contrast, fundamental KUs like Data Type and Loop, derived from the Java SE I exam, are more prevalent.
Summary RQ2: What are the most important KU features for predicting post-release defects? Method & Encapsulation, Inheritance, Exception Handling, Advanced Class Design, and Loop are the top five most important KU features for predicting post-release defects across all studied releases.
4.3 RQ3: Can KUM be made more accurate by combining its features with those from TM?
Motivation. This research question investigates whether the predictive performance of KUM for post-release defects can be improved by integrating them with traditional metrics. The underlying motivation is based on the hypothesis that combining diverse sets of features can provide a more comprehensive understanding of the factors influencing post-release defects. Such integration may enable the combined model to capture a wider range of predictive indicators, thereby enhancing its accuracy in identifying potential defects.
Approach. We build a combined model named KUM+TM for predicting post-release defects using all studied features of KUM and TM. Following Step 1 of RQ1 (Figure 3), we build KUM+TM models for all studied releases. We rank the studied prediction models (KUM, TM, PROD, PROC, OWN and KUM+TM) using SK-ESD (Ghotra et al., 2015; Mittas and Angelis, 2013; Tantithamthavorn et al., 2017). We further compare the performance of KUM+TM model with the performance of KUM and TM models for each of the studied releases. We also analyze feature importances of KUM+TM model using the same SHAP analysis method explained in Section 4.2.
Figure 6: The distribution of AUC of the combined KUM+TM and other studied models. The models are grouped based on their performance rankings determined by the Scott-Knott ESD (SK-ESD) method, where a lower SK-ESD rank indicates a better-performing model.